OpenStudy (anonymous):

What is the equation of the line that contains the point (–5, –1) and has a slope of 4? Write in slope-intercept form. Im sorry I really don't understand slope intercept

OpenStudy (anonymous):

The easiest way is to write it in point slope form first. To do this follow the form y-k = s(x-h) where k is your y coordinate point, and h is your x coordinate point and s is your slope. If we do this with your equation, what would we get?

OpenStudy (anonymous):

well the slope is 4 and -5,-1 are the coordinates all we do is plug them in right?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

I get it now! I didn't realize that's all you had to do is plug it in to the formula thanks guys:)

OpenStudy (anonymous):

That isn't all. That puts it in point-slope form.

OpenStudy (anonymous):

You want slope-intercept form, which is one extra step.

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

theres an extra step?

OpenStudy (anonymous):

Let's use an example, say my slope is 2, and I get a point of (1, 3). Point slope would be y-3 = 2(x-1) Slope intercept is the form y = mx + b If I make it easier to follow that form, y = 2(x-1) + 3 *Hint* What if we distribute the 2 ;)

OpenStudy (anonymous):

it would simplify the problem right?

OpenStudy (anonymous):

y = 2x - 2 + 3 which equals y = 2x + 1 Which form does y = 2x + 1 follow? Point slope, or your slope intercept?

OpenStudy (anonymous):

slope intercept because y=mx+b correct?

OpenStudy (anonymous):

Yes ma'am. :) Now just do the same with your problem. Put it into point slope form. And then get y on the left, and everything else on the right. And then distribute your slope and simplify. Problem = solved.

OpenStudy (anonymous):

thankyou:) no one has ever shown me that extra step!

OpenStudy (anonymous):

You are very welcome. Thank for your dedication to learning!

OpenStudy (anonymous):

well thankyou for making me understand as well as taking the time to explain it!